OpenStudy (anonymous):

K is the midpoint of line segment AB. if line segment AK = 4x + 3 and line segment KB = 7x-3, find the length of line segment AB A. 11 B. 2 C. 22 D. 4

OpenStudy (anonymous):

a answe is actually 11x

OpenStudy (anonymous):

Thank you very much! :)

OpenStudy (anonymous):

you are welcome

hero (hero):

The correct answer is B. 2 not a

hero (hero):

4x + 3 = 7x - 3 7x - 4x = 3 + 3 3x = 6 x = 6/3 x = 2

hero (hero):

Juanita's work is not correct

OpenStudy (anonymous):

Well I still appreciate that people are taking their time to try and help me out. But thank you Hero for showing me what was wrong! :)